OpenStudy (anonymous):

f(x)=x^2-4 and g(x)=2x+5 find f*g and g*f

OpenStudy (anonymous):

f*g(x)=f(g(x))=f(2x+5)=(2x+5)^2-4

OpenStudy (unklerhaukus):

\[*=\circ\] ?